# Reassessing Collaborative Writing Theories and Frameworks in the Age of LLMs: What Still Applies and What We Must Leave Behind

### v2

@@ 4.2.4. Coherence support is an ongoing challenge
While appropriate prompting could potentially be a solution for achieving lexical and syntactic alignment, semantic alignment (i.e. consistency in the flow and document structure) would require more intricate design intervention, since it is less about the LLM’s writing capabilities and more about the human-AI alignment. Visual representation of the semantic document structure as seen in VISAR (Zhang et al., 2023) serves as a great inspiration here. Providing a visual cue of the current document structure would not just allow the user to be in constant alignment with the AI, but would essentially force the LLMs to engage in some form of “planning”, which would not otherwise be technically needed (Elkins, 2024) (unless specified to do so, such as with Chain-of-Thought prompting). Ultimately, it would enable the user to have an understanding of “how their text works” (Rahhal et al., 2007), which will be valuable regardless of the mode of collaboration.
@@ 4.2.5. Writing is socially constructed by readers
If readers are a part of the writing, then it would mean that readers are inseparable from the LLM’s text generation process. A straightforward implementation would be to simply include the intended audience profile to the prompt, but we do not see this to be the solution. The fundamental implication here is that writers should always be conscious of the reader’s grammar, language and existing knowledge to maximize the ease of comprehension (Rahhal et al., 2007). One possible approach is to incorporate reader profile awareness into the revision process, where the AI co-writer is given the ability to assess the text as a pre-defined audience to check for any inconsistencies or lack of information from their perspective. The risk here would be excessive personalization, leading to bias amplification and social echo chambers, like we are starting to see in the adjacent field of human-LLM dialogue (Knoeferle et al., 2025).
@@ 4.2.6. Lack of clarity with authorship
The concept of authorship has been an ongoing debate in collaborative writing (Ede and Lunsford, 2001). While we excluded papers that primarily focused on authorship during our screening, it has still come up as an evident theme. For instance, Dhillon et al. (Dhillon et al., 2024) explored the effects of scaffolding during human-AI collaborative writing, and found that high scaffolding (i.e. next paragraph suggestion) resulted in higher writing quality and productivity, but also resulted in the decrease of text ownership and overall satisfaction. As more and more writing is conducted by AI, it is inevitable that our sense of ownership towards the text will decrease. Since our scope for this paper is the pragmatic support of task completion, we do not necessarily see this as a problem. Having said that, the concept of authorship is directly tied with the author’s accountability (Stokel-Walker, 2023), and given that only the human writers can be kept accountable for the consequences of any human-AI collaborative writing (Stokel-Walker, 2023), design considerations are necessary to clarify where the accountability lies.
Design implication #6: Design data structures and visual cues to distinguish human-written text from AI-written ones & Provide built-in verification methods for AI-generated text
We believe that keeping track of human contribution vs LLM contribution is a crucial aspect in future collaborative writing tools. Laban et al’s (Laban et al., 2024) InkSync proposes an insightful JSON data structure that can constantly keep track of text revisions as well as flagging for additional fact checking. Such data structures would allow for collaborative writing tools to visually distinguish human written text from AI written ones, enhancing the notion of accountability for the user. Their work also proposes the Warn-Verify-Audit framework, where new information introduced to the document by LLMs are flagged, and can be automatically searched online or sent to an external auditor for manual fact checking. We see similar interesting fact-checking approaches emerging (Liu et al., 2024; Amirizaniani et al., 2024; Dietz, 2024), all of which are human-in-the-loop approaches to maximize the strengths of manual oversight and LLM automation. While these studies are beyond the scope of this paper, they would inevitably need to be eventually built into human-AI collaborative writing tools.
@@ 4.2.7. Collaborative writing tools serve as external cognition
Fiore and Wiltshire (Fiore and Wiltshire, 2016) explored the various theories related to technology as an artifact of external cognition to make the claim that tools are inseparable from a team’s cognition. This allows us to circle back to the very reason why human-AI collaboration is beneficial in work-context writing tasks in the first place: the primary purpose of involving LLMs in collaborative writing should be to get the job done in the most effective manner, and the offloading of cognitive resources is a key criteria in achieving this goal. As we design AI co-writer tools, we must take great care in making sure that the tool is not resulting in unwanted cognitive load, bringing more disruption than improvement. Unnecessary cognitive load can be witnessed at various moments throughout the writing process. For example, Dhillon et al. (Dhillon et al., 2024) found that low-level scaffolding (i.e. next-sentence suggestions) did not enhance writing productivity, due to its intended assistance being too “rudimentary and disruptive” for the user’s writing process. Instead, high-level scaffolding (i.e. next-paragraph suggestions) resulted in significant improvements in the writing quality, due to the effective reduction of cognitive load. Larsen-Ledet et al. (Larsen-Ledet et al., 2020) shed light from a socio-technical perspective, finding that writers often switch between multiple writing tools, not just for functional reasons, but to get away from the collaborative mode to “sit for myself.” That is, writers sometimes prefer to stay away from collaboration to stay focused.
Traditionally, cognitive load has been measured in HCI through post-task questionnaires, such as with the NASA task load index (Hart and Staveland, 1988), although recently the validated of this for HCI has been questioned (Babaei et al., 2025). Emerging frameworks such as the Overall Writing Effectiveness framework (Wilbers et al., 2024) could also serve as an alternative metric in surfacing the impacts that LLMs bring to collaborative writing.
@@
When working collaboratively, it is not just the task that becomes the source of cognitive effort, but also the process itself, implying that the AI co-writer’s role is not limited to producing high quality text, but also to make sure that the process itself is not causing discomfort for the user. This is inspired greatly by the way Fiore and Wiltshire (Fiore and Wiltshire, 2016) emphasize the importance of distinguishing teamwork (behaviors required for working together with others) and taskwork (functions required for getting the task done). Tools support teamwork through the visualization of group awareness, team progress and facilitating communication, while tools also support taskwork through the direct assistance of the task (e.g. writing or decision making). In other words, collaborative writing tools can help to offload cognitive load not just through the direct contribution to the writing as a co-writer, but also by streamlining the collaboration process itself, both for human-AI teamwork as well as human-human teamwork.
@@ 5. Discussion
While the role of humans undeniably changes from a hands-on writer to more of an editorial decision maker (Buschek, 2024), we believe that the cognitive process of writing will not and should not change — at least for good writing. This is to make sure that the writer continues to stay accountable for the writing they produce (which the AI cannot). Because of this, the non-linear, interchangeable nature of writing that studies based on the Cognitive Process Theory of Writing (Flower and Hayes, 1981) will stay as a core foundation of future collaborative writing tool designs. However, because the speed and ease of translating ideas to words will increase dramatically, we believe the writing process to become closer to an empirical “trial and error” process analogous to prototyping, resulting in less need for fully fledged planning, and more support needed for rigorous revision methodologies. In particular, tools will need to equip writers with enhanced understanding and terminologies for different revision types, to draw more granular boundaries for the AI support.
@@
This shift in turn would further elevate the importance of coherence support in human-AI collaborative writing, which has been an ongoing challenge in this field to this day. Given the unprecedented semantic classification capabilities of LLMs, we are optimistic about the impact that LLMs can bring to the alignment of document structures through real-time visualization, which can positively impact not just the human-AI collaborative writing but human-human writing as well.
@@
Since we take the stance that AIs are no more than a tool and humans should always remain in control, teamwork-related factors such as group awareness, consensus building and authorship, which have been much discussed in collaborative writing thus far, will not apply in the human-AI paradigm, except between multiple human co-authors. Therefore, when it comes to the writing strategies, it seems inevitable that most human-AI collaborative writing will end up as single-author writing, or will be conducted in the hierarchical mode where the AI co-writer will primarily translate ideas to words instantaneously under human supervision, and the humans will primarily take an editorial role. The more equally distributed forms of collaborative writing, known as reactive writing, parallel writing or the dialogic mode, will remain a strategy unique to human collaboration. However, given that the reduction of cognitive load results not just from the direct support of text production but in streamlining the collaboration process as well, it is an equally important role of the AI co-writer to remove any unnecessary frictions in the collaboration process itself, particularly when there are more than one human writers involved.
@@ 6. Limitations
Finally, we would like to acknowledge some general limitations of our literature analysis approach.
First, we chose to look at Web of Science and the ACM Digital Library, but it is possible that we are missing collaborative writing theories or frameworks that were not covered by these two databases. We are, however, confident that we have sufficiently covered at least the most influential ones commonly found in past literature.
@@ 7. Conclusion
In this paper, we have argued for the need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm, as LLMs increasingly become our daily co-writers in a uniquely asymmetrical relationship unlike any we have previously encountered. We conducted a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the human-AI paradigm, and drew seven insights and implications that can be applied to the design of human-AI writing tools. We found that, as LLMs nudge the writing process more towards an empirical “trial and error” process analogous to prototyping, the non-linear cognitive process of writing will stay the same, but more attention will be required for rigorous revision methodologies. In turn, this may reduce the focus on the planning aspects of collaborative writing, further increasing the importance of coherence support between human and AI writing. We believe that the LLM’s unprecedented semantic capabilities can bring a break through to this ongoing challenge, through real time visualizations of the document’s semantic structure. Teamwork-related factors such as group awareness, consensus building and authorship, which have been central discussions in human-human collaborative writing, will not apply to the human-AI paradigm, resulting in the majority of human-AI collaborative writing being categorized in the single-author writing strategy. Finally, writing will continue to be socially constructed by readers, but this is dependent on the fact that humans will continue to be the primary readers of the documents we create – which may not necessarily be the case in the near future. Beyond theoretical insights, we provide practical design implications and example interface sketches to guide real-world application. We hope that our work will bring clarity and theoretical grounding to the interaction designs of human-AI collaborative writing.
@@ Acknowledgements.
The interface designs in this paper used Google’s Material Icons444See https://fonts.google.com/icons provided under the Apache License, Version 2.0.References
@@
- Babaei et al. (2025) Ebrahim Babaei, Tilman Dingler, Benjamin Tag, and Eduardo Velloso. 2025. Should we use the NASA-TLX in HCI? A review of theoretical and methodological issues around Mental Workload Measurement. International journal of human-computer studies 201, 103515 (June 2025), 103515. doi:10.1016/j.ijhcs.2025.103515
@@
- Hart and Staveland (1988) Sandra G Hart and Lowell E Staveland. 1988. Development of NASA-TLX (task load index): Results of empirical and theoretical research. In Advances in Psychology. Advances in psychology, Vol. 52. Elsevier, 139–183.
@@
- Wang et al. (2020) Dakuo Wang, Elizabeth Churchill, Pattie Maes, Xiangmin Fan, Ben Shneiderman, Yuanchun Shi, and Qianying Wang. 2020. From human-human collaboration to human-AI collaboration. In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA.

### v3

@@ 4.2.4. Coherence support is an ongoing challenge
While appropriate prompting could potentially be a solution for achieving lexical and syntactic alignment, semantic alignment (i.e. consistency in the flow and document structure) would require more intricate design intervention, since it is less about the LLM’s writing capabilities and more about the human-AI alignment. Visual representation of the semantic document structure as seen in VISAR (Zhang et al., 2023) serves as a great inspiration here. Providing a visual cue of the current document structure would not just allow the user to be in constant alignment with the AI, but would also involve the LLMs to engage in a form of “planning”, which they do not require otherwise (Elkins, 2024) (unless specified to do so, such as with Chain-of-Thought prompting). Ultimately, it would enable the user to have an understanding of “how their text works” (Rahhal et al., 2007), which will be valuable regardless of the mode of collaboration.
@@ 4.2.5. Writing is socially constructed by readers
If readers are a part of the writing, then it would mean that readers are inseparable from the LLM’s text generation process. A straightforward implementation would be to simply include the intended audience profile to the prompt, but we do not see this to be a solution. The fundamental implication here is that writers should always be conscious of the reader’s grammar, language and existing knowledge to maximize the ease of comprehension (Rahhal et al., 2007). One possible approach is to incorporate reader profile awareness into the revision process, where the AI co-writer is given the ability to assess the text as a pre-defined audience to check for any inconsistencies or lack of information from their perspective. The risk here would be excessive personalization, leading to bias amplification and social echo chambers, like we are starting to see in the adjacent field of human-LLM dialogue (Knoeferle et al., 2025).
@@ 4.2.6. Lack of clarity with authorship
The concept of authorship has been an ongoing debate in collaborative writing (Ede and Lunsford, 2001). While we excluded papers that primarily focused on authorship during our screening, it has still come up as an evident theme. For instance, Dhillon et al. (Dhillon et al., 2024) explored the effects of scaffolding during human-AI collaborative writing, and found that high scaffolding (i.e. suggestions for entire paragraphs) resulted in higher writing quality and productivity, but also resulted in the decrease of text ownership. As more and more writing is conducted by AI, it seems inevitable that our sense of ownership towards the text will decrease(Kosmyna et al., 2025). Since our scope for this paper is the pragmatic support of task completion, we do not necessarily see this as a problem. Having said that, the concept of authorship is directly tied with the author’s accountability (Stokel-Walker, 2023), and given that only the human writers can be kept accountable for the consequences of any human-AI collaborative writing (Stokel-Walker, 2023), design considerations to clarify where the accountability lies seems necessary.
Design implication #6: Design data structures and visual cues to distinguish human-written text from AI-written ones
We believe that keeping track of human contribution vs LLM contribution is a crucial aspect in future collaborative writing tools. Laban et al’s (Laban et al., 2024) InkSync proposes an insightful JSON data structure that can constantly keep track of text revisions as well as flagging for additional fact checking. Such data structures would allow for collaborative writing tools to visually distinguish human written text from AI written ones, enhancing the notion of accountability for the user. In Kosmyna et al. (Kosmyna et al., 2025)’s study on essay writing with ChatGPT, they also touch on how lower ownership of leads to the “diminished sense of cognitive agency”. That is, when we feel that a written text is AI’s and not ours, we will feel less control over the product, leading to decrease in error monitoring. Perhaps then it is important to go beyond visualization of writers, and to provide interactions that will encourage human writers to increase their sense of ownership over the text.
@@ 4.2.7. Collaborative writing tools serve as external cognition
Fiore and Wiltshire (Fiore and Wiltshire, 2016) explored the various theories related to technology as an artifact of external cognition to make the claim that tools are inseparable from a team’s cognition. This allows us to circle back to the very reason why human-AI collaboration is beneficial in work-context writing tasks in the first place: the primary purpose of involving LLMs in collaborative writing should be to get the job done in the most effective manner, and the offloading of cognitive resources is a key criteria in achieving this goal. As we design AI co-writer tools, we must take great care in making sure that the tool is not resulting in additional cognitive load, bringing more disruption than improvement. Undesired cognitive load can be witnessed at various moments throughout the writing process, not just in the writing. For example, Dhillon et al. (Dhillon et al., 2024) found that low-level scaffolding (i.e. next-sentence suggestions) did not enhance writing productivity, due to its intended assistance being too “rudimentary and disruptive” for the user’s writing process. Instead, high-level scaffolding (i.e. next-paragraph suggestions) resulted in significant improvements in the writing quality, due to the effective reduction of cognitive load. Larsen-Ledet et al. (Larsen-Ledet et al., 2020) shed light from a socio-technical perspective, finding that writers often switch between multiple writing tools, not just for functional reasons, but to get away from the collaborative mode to “sit for myself.” That is, writers sometimes prefer to stay away from collaboration to stay focused.
@@
When working collaboratively, it is not just the task that becomes the source of cognitive effort, but also the process itself, implying that the AI co-writer’s role is not limited to producing high quality text, but also to make sure that the process itself is not causing discomfort for the user. This is inspired greatly by the way Fiore and Wiltshire (Fiore and Wiltshire, 2016) emphasize the importance of distinguishing teamwork (behaviors required for working together with others) and taskwork (functions required for getting the task done). Tools can support teamwork through the visualization of group awareness, team progress and facilitating communication, while tools can also support taskwork through the direct assistance of the task (e.g. writing or decision making). In other words, collaborative writing tools can help to offload cognitive load not just through the direct contribution to the writing as a co-writer, but also by streamlining the collaboration process itself, both for human-AI teamwork as well as human-human teamwork.
@@ 5. Discussion
5.1. Applying existing human-human collaborative writing theories/frameworks to the human-AI paradigm
@@
While the role of humans will change from a hands-on writer to more of an editorial decision maker (Buschek, 2024), we believe that the cognitive process of writing will not and should not change — at least for good writing. This is to make sure that the writer continues to stay accountable for the writing they produce (which the AI cannot) (Stokel-Walker, 2023). Because of this, the non-linear, interchangeable nature of writing based on the Cognitive Process Theory of Writing (Flower and Hayes, 1981) should stay as a core foundation of future collaborative writing tool designs. However, as more humans choose to delegate the actual writing to the essentially effortless AI generation, the legacy planning/translating/reviewing process shifts to a planning/waiting/reviewing process, thereby making the human writer’s cognitive process discontinuous due to the waiting that occurs in between. It becomes crucial then for the collaborative writing tool to provide support in streamlining the otherwise broken cognitive process of the human writer. For this we suggest a “prototyping” approach, where the tool works to shorten the waiting stage as well as accelerate the overall cognitive process, to allow human writers to test new ideas at a similar cognitive speed as when they are writing themselves. For instance, tools could enable quick iterations of the cognitive process by starting with digestible, smaller chunks of text, and gradually moving on to a fully fleshed-out document, much like evolving prototypes from low-fidelity to high-fidelity. Consequently, more rigor will be required for novel revision methodologies and interactions to allow for human writers to quickly read, review and edit AI generated text. Tools will need to equip writers with enhanced understanding and terminologies for various revision types, to set more granular boundaries for the desired AI support and to draw out the best of human and AI performance.
@@
As AI takes on more of the content writing, the ongoing challenge of coherence support will be of further importance. Given that LLMs provide a level of semantic generalization and abstraction not achievable with prior AI models, we are optimistic about the impact that LLMs can bring to the alignment of document structures, such as through real-time visualization. These interventions will positively impact not just the human-AI collaborative writing but human-human writing as well.
@@
Since we take the stance that AIs are no more than a tool and humans should always remain in control, teamwork-related factors such as group awareness, consensus building and authorship, which have been much discussed in collaborative writing studies thus far, will not apply in the human-AI paradigm, except between multiple human co-authors. Therefore, when it comes to the writing strategies, it seems inevitable that most human-AI collaborative writing will end up as single-author writing, or will be conducted in the hierarchical mode where the AI co-writer will primarily translate ideas to words instantaneously under human supervision, and the humans will primarily take an editorial role. The more equally distributed forms of collaborative writing, known as reactive writing, parallel writing or the dialogic mode, will remain a strategy unique to human collaboration. However, given that the reduction of cognitive load derives not just from the direct support of text production but from the streamlining of the overall collaboration process, it is equally important that the AI co-writer works to remove any unnecessary frictions in the collaboration process itself, particularly when there are more than one human writers involved.
@@
Much like how word-processors had a profound impact on the way that we wrote with pen and paper, LLMs seem to be fundamentally shifting the way that we approach writing. While there are certainly signs of de-skilling, as HCI researchers we are also curious to see what synergistic effects arise from this shift, resulting in novel mindsets and methodologies of writing. As Brinkmann et al. (Brinkmann et al., 2023) argue through the lens of “machine cultures”, our culture of writing is already being reshaped by our interactions with LLMs.
5.2. Prompting as a form of writing
In Figure 1, we assumed that, when writing with AI, prompting becomes the key component of the planning process. When prompting LLMs, we consider not just the content of the writing but the audience, length, structure, tone, and internal knowledge that we posses (often referred to as “context” in prompt engineering) that is required for the writing. This is much in line with how Flower and Hayes describe their model: they use the term planning in a “much broader sense”, where writers form abstract “representations” of ideas and goals, often involving retrieval from long-term memory (Flower and Hayes, 1981). The way they explain the iterative nature of goal-setting is also highly analogous to the way LLMs are used today, refining instructions through multiple trial and errors (Arawjo et al., 2023; Liu et al., 2022).
However, it is important to note that prompting itself is also a form of writing. That is, the act of writing a prompt is also the very result of the writer’s cognitive writing process. In this scenario, the user would first engage in the planning/translating/reviewing to write a prompt, which would then be fed to an LLM for the user to engage this time in a planning/waiting/reviewing process. Further, if the writer chooses to use meta-prompting (Reynolds and McDonell, 2021) (using LLMs to create and refine prompts), then they go through the planning/translating/reviewing process to write a prompt that retrieves a prompt which they would revise and iterate through the planning/waiting/reviewing process, then apply the generated prompt to the LLM to generate the final outcome which again would be revised through the planning/waiting/reviewing process, and so forth. In either case, when viewed from the lens of this study, prompt writing is distinctly different from most forms of writing in the sense that it is a uniquely reader specific form of writing –in this case the reader of course being the LLM. Grammar or spelling errors are not rigorously checked, since they do not matter as long as the instruction gets through. Additional contexts are also often added without much thought, regardless of knowing whether the outcome would be improved from doing so(Liu et al., 2022). Prompt engineering or defense techniques (Liu et al., 2023) (such as the ”sandwich prevention”, which refers to repeating the certain instructions at the beginning and the end of a prompt) are often highly unnatural forms of writing. Despite these peculiarities as a form of writing, it is also interesting that the concept of authorship seems to apply to prompts: platforms such as PromptBase444See https://promptbase.com/ allow users to sell and buy effective prompts. As such, prompting itself can be seen as the act of writing, with distinct characteristics. In this sense, when humans co-write with AI, the process begins with a legacy human mode of writing, resulting in a mix of both approaches that we outlined in Figure 11. How does this affect the overall process of writing with AI? Is it necessary to intentionally separate the initial prompting (i.e. the legacy planning/translating/reviewing process) from the human-AI mode of writing (i.e. the planning/waiting/reviewing process) that follows? These are still early questions in trying to understand how LLMs are affecting the way that we write, and we welcome more researchers to explore this space.
@@ 6. Limitations
Finally, we acknowledge some general limitations of our literature analysis approach.
First, we chose to look at Web of Science and the ACM Digital Library, but it is possible that we are missing collaborative writing theories or frameworks that were not covered by these two databases. For example, among the 6 articles that we added through backward referencing, 4 were found to be in the ProQuest database (the remaining 2 were the cognitive process theories by Flower and Hayes (Flower and Hayes, 1981; Hayes, 1996), which are fundamental to this study but are technically not collaborative writing). We are, however, confident that we have sufficiently covered at least the most influential ones commonly found in past literature.
@@ 7. Conclusion
In this paper, we have argued for the need to reassess foundational theories and frameworks of collaborative writing to adapt to the human-AI paradigm, as LLMs increasingly become our daily co-writers. We conducted a critical review of existing theories and frameworks on human-human collaborative writing to assess their relevance to the human-AI paradigm, and drew seven insights and implications that can be applied to the design of human-AI writing tools. We found that, as we delegate more writing to AI, our cognitive process shifts from the traditional planning/translating/reviewing process to a planning/waiting/reviewing process, breaking the process due to the waiting that occurs in between. To ensure that our cognitive process remains intact, we suggest a “prototyping” approach, in which the tool allows for quick iterations of the cognitive process by starting with smaller or summarized chunks of text, and gradually moving on to a fully fleshed document. For this we would need to come up with novel revision methodologies and interactions. In turn, this may reduce the focus on the planning aspects of collaborative writing, further increasing the importance of coherence support between human and AI writing. We believe that the LLM’s strong semantic capabilities can provide analyses that were previously difficult with traditional methods, such as real time visualizations of a document’s semantic structure. Teamwork-related factors such as group awareness, consensus building and authorship, which have been central discussions in human-human collaborative writing, will not apply to the human-AI paradigm, resulting in the majority of human-AI collaborative writing being categorized in the single-author writing strategy. Finally, writing will continue to be socially constructed by readers, but this is dependent on the fact that humans will continue to be the primary readers of the documents we create – which may not necessarily be the case in the near future. Beyond theoretical insights, we provide practical design implications and example interface sketches to guide real-world application. We hope that our work will bring clarity and theoretical grounding to the interaction designs of human-AI collaborative writing.
@@ Acknowledgements.
The interface designs in this paper used Google’s Material Icons555See https://fonts.google.com/icons provided under the Apache License, Version 2.0.References
@@
- Arawjo et al. (2023) Ian Arawjo, Priyan Vaithilingam, Martin Wattenberg, and Elena Glassman. 2023. ChainForge: An open-source visual programming environment for prompt engineering. In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, New York, NY, USA.
- Brinkmann et al. (2023) Levin Brinkmann, Fabian Baumann, Jean-François Bonnefon, Maxime Derex, Thomas F Müller, Anne-Marie Nussberger, Agnieszka Czaplicka, Alberto Acerbi, Thomas L Griffiths, Joseph Henrich, Joel Z Leibo, Richard McElreath, Pierre-Yves Oudeyer, Jonathan Stray, and Iyad Rahwan. 2023. Machine culture. Nat. Hum. Behav. 7, 11 (Nov. 2023), 1855–1868.
@@
- Kosmyna et al. (2025) Nataliya Kosmyna, Eugene Hauptmann, Ye Tong Yuan, Jessica Situ, Xian-Hao Liao, Ashly Vivian Beresnitzky, Iris Braunstein, and Pattie Maes. 2025. Your brain on ChatGPT: Accumulation of cognitive debt when using an AI assistant for essay writing task. arXiv [cs.AI] (June 2025).
@@
- Lee et al. (2025) Hao-Ping (hank) Lee, Advait Sarkar, Lev Tankelevitch, Ian Drosos, Sean Rintel, Richard Banks, and Nicholas Wilson. 2025. The impact of generative AI on critical thinking: Self-reported reductions in cognitive effort and confidence effects from a survey of knowledge workers. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA, 1–22.
@@
- Liu et al. (2022) Jiachang Liu, Dinghan Shen, Yizhe Zhang, Bill Dolan, Lawrence Carin, and Weizhu Chen. 2022. What makes good in-context examples for GPT-3?. In Proceedings of Deep Learning Inside Out (DeeLIO 2022): The 3rd Workshop on Knowledge Extraction and Integration for Deep Learning Architectures. Association for Computational Linguistics, Stroudsburg, PA, USA, 100–114.
- Liu et al. (2023) Yupei Liu, Yuqi Jia, Runpeng Geng, Jinyuan Jia, and Neil Zhenqiang Gong. 2023. Formalizing and benchmarking prompt injection attacks and defenses. arXiv [cs.CR] (Oct. 2023).
@@
- Ma et al. (2024) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma. 2024. Towards human-AI deliberation: Design and evaluation of LLM-empowered deliberative AI for AI-assisted decision-making. arXiv [cs.HC] (March 2024).
@@
- Naikar et al. (2025) Neelam Naikar, Robert Hoffman, Emilie M Roth, Gary Klein, Laura G Militello, and Cindy Dominguez. 2025. Should we Make AI More Tool-like or Teammate-Like? J. Cogn. Eng. Decis. Mak. 15553434251346904 (June 2025).
@@
- Reynolds and McDonell (2021) Laria Reynolds and Kyle McDonell. 2021. Prompt programming for large language models: Beyond the few-shot paradigm. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, New York, NY, USA.
@@
- Zhu et al. (2024) Tiffany Zhu, Iain Weissburg, Kexun Zhang, and William Yang Wang. 2024. Human bias in the face of AI: The role of human judgement in AI generated text evaluation. arXiv [cs.CL] (Sept. 2024).